# Dark Matter Physics in General NMSSM

### v1

@@ 1 Introduction
WIMPs can naturally arise in the supersymmetric extension of the SM in particle physics. Furthermore, supersymmetric theories possess numerous intriguing features, such as the absence of quadratic divergences, gauge coupling unification, and distinctive signatures of supersymmetric particles produced in present and future colliders. In the framework of the Minimal Supersymmetric Standard Model (MSSM), recognized as one of the most economical supersymmetric theories, the neutral electroweakino sector encompasses the mixing of Bino (denoted as $\tilde{B}$), Wino (referred to as $\tilde{W}$), and neutral Higgsinos (designated as $\tilde{H}_{u}$ and $\tilde{H}_{d}$), giving rise to the formation of four distinct neutralinos [16, 17, 18]. Among these, the lightest mass eigenstate, denoted as $\tilde{\chi}_{1}^{0}$, usually corresponds to the lightest supersymmetric particle (LSP) and acts as a feasible candidate of DM. It prefers to be $\tilde{B}$-dominated and co-annihilate with the $\tilde{W}$-like electroweakinos to achieve the measured relic abundance. As indicated by the research in Ref. [19], the LZ experiment alone imposes a lower bound on the Higgsino mass, $\mu\gtrsim 380$ GeV, in the absence of strong cancellations between different contributions to the SI scattering. Although such a large $\mu$ may be naturally generated by the well-known Giudice–Masiero mechanism in the gravity-mediated supersymmetry breaking scenario [20], it leads to severe fine-tuning problems in precisely predicting the mass of the $Z$ boson, after considering the Large Hadron Collider (LHC) Higgs discovery and non-observations of any DM and supersymmetric particle signals in experiments when the theory runs down from an infrared high-energy scale to the electroweak scale [21, 22, 23].
@@
The dilemma faced by the MSSM inspired us to extend this model. The Next-to-Minimal Supersymmetric Standard Model with a $\mathbb{Z}_{3}$ symmetry ($\mathbb{Z}_{3}$-NMSSM) introduces a gauge singlet superfield $\hat{S}$ on top of the MSSM [24, 25]. In this framework, the $\mu$-parameter of the MSSM is dynamically generated once the scalar component of the superfield $\hat{S}$ acquires a vacuum expectation value (vev) below approximately 1 TeV. This places it naturally at the electroweak scale. The neutralino sector includes an additional fermionic partner of $\hat{S}$, known as the Singlino. Both the Bino-dominated (which is the case in most physical scenarios) and Singlino-dominated neutralinos are viable DM candidates [26, 27, 28, 5, 29, 30, 31, 32, 33, 34, 15, 35, 36, 37, 38, 39, 40]. Primarily, the Bino-dominated DM candidate differs from the MSSM given the former’s ability to co-annihilate with a Singlino-dominated neutralino in achieving the measured abundance. Nonetheless, this scenario is confined to an exceedingly limited parameter space defined by specific conditions: $|2\kappa\mu/\lambda|\simeq|M_{1}|$, moderately large $\lambda$ and $\kappa$, and $\mu$ values exceeding 300 GeV [5, 15]. This space predicts a significant mixing of the SM Higgs field with the singlet Higgs field and is thus disfavored by the Higgs property measurements at the LHC. The characteristics of Singlino-dominated DM are contingent upon the values of $\lambda$, $\tan\beta$, $m_{\tilde{\chi}_{1}^{0}}$, and $\mu_{\rm eff}\equiv\lambda v_{s}/\sqrt{2}$  [41], where $\lambda$ denotes the singlet–doublet Higgs Yukawa coupling within the superpotential and $m_{\tilde{\chi}_{1}^{0}}\simeq 2\kappa\mu/\lambda$ requires $12|\kappa|/\lambda<$ 1. Considering that both SI and SD DM-nucleon scattering cross sections scale with $\lambda^{4}$ in the regime of heavy singlet Higgs bosons and consequently, the LZ experiment generally sets an upper limit of $\lambda\lesssim 0.1$, the DM is likely to achieve the observed abundance through one of two main mechanisms [29]: co-annihilating with Higgsino-dominated electroweakinos or undergoing resonant annihilations facilitated by singlet-dominated CP-even or CP-odd Higgs bosons. The former scenario is only viable within a narrow parameter space characterized by $2|\kappa|/\lambda\simeq 1$, $\lambda<0.1$, and $\mu<400$ GeV, where $\kappa$ represents the self-coupling coefficient of the singlet Higgs field [41]. To achieve the measured density, the latter scenario requires $2|m_{\tilde{\chi}_{1}^{0}}|$ to be close to the scalar mass. Thus, all these scenarios correspond to an exceptionally constrained parameter space, resulting in a significant suppression of Bayesian evidence [^1] Bayesian evidence is a fundamental concept in Bayesian statistics [42]. It is typically denoted as $Z(D|M)\equiv\int{P(D|O(M,\Theta))P(\Theta|M)\prod d\Theta_{i}}$, where $P(\Theta|M)$ represents the prior probability density function of input parameters $\Theta=(\Theta_{1},\Theta_{2},\cdots)$ in model $M$, and $P(D|O(M,\Theta))\equiv\mathcal{L}(\Theta)$ is the likelihood function for observed values $O$. The Bayesian evidence considers both the theoretical predictions $O(M,\Theta)$ and the experimental data $D$, and its computation involves marginalizing the model’s parameters, which requires integrating the likelihood function and prior probability over all possible parameter values. In Bayesian statistics, evidence is used to compare different models and determine which one is more plausible for explaining the observed data. While a higher value of $Z$ suggests that the respective model is more likely to agree with the data, an extremely small $Z$ means that the theory requires its parameters to be fine-tuned for the same..
@@ 2 Theoretical preliminaries
$W_{\rm GNMSSM}=W_{\text{Yukawa}}+\lambda\hat{S}\hat{H_{u}}\cdot\hat{H_{d}}+\frac{\kappa}{3}\hat{S}^{3}+\mu\hat{H_{u}}\cdot\hat{H_{d}}+\xi\hat{S}+\frac{1}{2}\mu^{\prime}\hat{S}^{2},$ (2.1)
@@ 2.1 Higgs sector of GNMSSM
$-\mathcal{L}_{soft}=$ $\Bigg{[}\lambda A_{\lambda}SH_{u}\cdot H_{d}+\frac{1}{3}\kappa A_{\kappa}S^{3}+m_{3}^{2}H_{u}\cdot H_{d}+\frac{1}{2}{m_{S}^{\prime}}^{2}S^{2}+h.c.\Bigg{]}$ (2.2)
@@
with $v=\sqrt{v_{u}^{2}+v_{d}^{2}}\simeq 246~{}\mathrm{GeV}$, and the three masses are determined in solving the conditional equations to minimize the scalar potential. The Higgs sector is then described by ten free parameters:
$\tan\beta,~{}\lambda,~{}\kappa,~{}v_{s},~{}A_{\lambda},~{}A_{\kappa},~{}\mu,~{}\mu^{\prime},~{}m_{3}^{2},~{}{m_{\text{S}}^{\prime}}^{2},$ (2.4)
@@
${\cal M}^{2}_{S,11}$ $=$ $\frac{\lambda v_{s}(\sqrt{2}A_{\lambda}+\kappa v_{s}+\sqrt{2}\mu^{\prime})+2m_{3}^{2}}{\sin 2\beta}+\frac{1}{2}(2m_{Z}^{2}-\lambda^{2}v^{2})\sin^{2}2\beta,$
@@
${\cal M}^{2}_{S,33}$ $=$ $\frac{(A_{\lambda}+\mu^{\prime})\sin 2\beta}{2\sqrt{2}v_{s}}\lambda v^{2}+\frac{\kappa v_{s}}{\sqrt{2}}(A_{\kappa}+2\sqrt{2}\kappa v_{s}+3\mu^{\prime})-\frac{\mu}{\sqrt{2}v_{s}}\lambda v^{2}.\quad$ (2.5)
@@
$\kappa A_{\kappa}$ $=$ $\frac{\sqrt{2}m^{2}_{B}}{v_{s}}+\frac{\lambda\mu v^{2}}{v_{s}^{2}}-\frac{\lambda(A_{\lambda}+\mu^{\prime})v^{2}\sin{2\beta}}{2v_{s}^{2}}-2\sqrt{2}\kappa^{2}v_{s}-3\kappa\mu^{\prime},$
$m_{S}^{\prime 2}$ $=$ $-\frac{1}{2}\left[m^{2}_{C}+\frac{\lambda\mu v^{2}}{\sqrt{2}v_{s}}+\frac{\kappa v_{s}(3A_{\kappa}+\mu^{\prime})}{\sqrt{2}}-\frac{\lambda(A_{\lambda}+2\sqrt{2}\kappa v_{s}+\mu^{\prime})v^{2}\sin{2\beta}}{2\sqrt{2}v_{s}}\right],$ (2.9)
@@ 2.3 DM physics of GNMSSM
In the GNMSSM, the following interactions are pertinent to the DM physics:
@@
$C_{h_{i}h_{s}h_{s}}$ $\simeq$ $\lambda v\,V_{h_{i}}^{\rm SM}V_{h_{s}}^{\rm S}V_{h_{s}}^{\rm S}(\lambda-\kappa\sin 2\beta)-\lambda\kappa vV_{h_{i}}^{\rm NSM}V_{h_{s}}^{\rm S}V_{h_{s}}^{\rm S}\cos 2\beta$ (2.19)
$+\sqrt{2}V_{h_{i}}^{\rm S}V_{h_{s}}^{\rm S}V_{h_{s}}^{\rm S}(3\sqrt{2}\kappa^{2}v_{s}+\kappa A_{\kappa}+3\kappa\mu^{\prime}),$
$C_{h_{i}A_{s}A_{s}}$ $\simeq$ $\lambda v\,V_{h_{i}}^{\rm SM}(\lambda+\kappa\sin 2\beta)+\lambda\kappa vV_{h_{i}}^{\rm NSM}\cos 2\beta$ (2.20)
$+\sqrt{2}V_{h_{i}}^{S}(\sqrt{2}\kappa^{2}v_{s}-\kappa A_{\kappa}+\kappa\mu^{\prime}).$
In the small $\lambda$ case, we conclude that
$C_{\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}h_{s}}\simeq-\sqrt{2}\kappa,\quad C_{h_{s}h_{s}h_{s}}\simeq\sqrt{2}(3\sqrt{2}\kappa^{2}v_{s}+\kappa A_{\kappa}+3\kappa\mu^{\prime}),$
$C_{\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}A_{s}}\simeq-\sqrt{2}\kappa,\quad C_{h_{s}A_{s}A_{s}}\simeq\sqrt{2}(\sqrt{2}\kappa^{2}v_{s}-\kappa A_{\kappa}+\kappa\mu^{\prime}),$ (2.21)
where $A_{\kappa}$ and $\mu^{\prime}$ are determined by $\lambda$, $\kappa$, $\tan\beta$, $v_{s}$, $A_{\lambda}$, $\mu_{\rm tot}$, $m_{N}$, and $m_{B}$ from Eq. (2.9), and the other interactions, such as $C_{\tilde{\chi}^{0}_{1}\tilde{\chi}^{0}_{1}h}$, $C_{\tilde{\chi}^{0}_{1}\tilde{\chi}^{0}_{1}Z}$, $C_{hh_{s}h_{s}}$, and $C_{hA_{s}A_{s}}$, are suppressed by $\lambda$.
@@ 2.3.1 DM relic density
    This formula together with Eq. (2.21) indicate that $\left\langle\sigma v\right\rangle_{x_{F}}^{h_{s}h_{s}}$ depends not only on $\kappa$, $m_{h_{s}}$, and $m_{\tilde{\chi}_{1}^{0}}$, but also on $\lambda$, $\tan\beta$, $v_{s}$, $\mu_{\rm tot}$, and $A_{\lambda}$ via $C_{h_{s}h_{s}h_{s}}$. In particular, $\kappa$ plays a critical role in determining the magnitude of $\left\langle\sigma v\right\rangle_{x_{F}}^{h_{s}h_{s}}$, since most of its dominant contributions are proportional to $\kappa^{4}$. If $C_{h_{s}h_{s}h_{s}}$ is sufficiently small so that the s-channel contribution can be neglected, we have
@@
    respectively, assuming that $h_{s}$ is not exceptionally light to forbid $A_{s}$-mediated resonant annihilation. Furthermore, from the formula of $C_{h_{s}A_{s}A_{s}}$ in Eq. (2.21), one can infer that if $|\sqrt{2}\kappa v_{s}-A_{\kappa}+\mu^{\prime}|\ll(4m_{\tilde{\chi}_{1}^{0}}^{2}-m_{A_{s}}^{2})/m_{\tilde{\chi}_{1}^{0}}$, $|{\cal{A}}_{t}|$ is much larger than $|{\cal{A}}_{s}|$, reflecting that the $t$-channel contribution prevails. In this case, $\left\langle\sigma v\right\rangle_{x_{F}}^{h_{s}A_{s}}$ is simplified as
@@
    if $m_{\tilde{\chi}_{1}^{0}}$ is much larger than $m_{A_{s}}$ and $m_{h_{s}}$, and

### v2

@@ 1 Introduction
WIMPs can naturally arise in the supersymmetric extension of the SM in particle physics. Furthermore, supersymmetric theories possess numerous intriguing features, such as the absence of quadratic divergences, gauge coupling unification, and distinctive signatures of supersymmetric particles produced in present and future colliders. In the framework of the Minimal Supersymmetric Standard Model (MSSM), recognized as one of the most economical supersymmetric theories, the neutral electroweakino sector encompasses the mixing of Bino (denoted as $\tilde{B}$), Wino (referred to as $\tilde{W}$), and neutral Higgsinos (designated as $\tilde{H}_{u}$ and $\tilde{H}_{d}$), giving rise to the formation of four distinct neutralinos [16, 17, 18]. Among these, the lightest mass eigenstate, denoted as $\tilde{\chi}_{1}^{0}$, usually corresponds to the lightest supersymmetric particle (LSP) and acts as a feasible candidate of DM. It prefers to be $\tilde{B}$-dominated and co-annihilate with the $\tilde{W}$-like electroweakinos to achieve the measured relic abundance. As indicated by the research in Ref. [19], the LZ experiment alone imposes a lower bound on the Higgsino mass, $\mu\gtrsim 380$ GeV, in the absence of strong cancellations between different contributions to the SI scattering. This bound can be enhanced by several tens of GeV after incorporating the radiative correction to the scattering [20, 21]. Although such a large $\mu$ may be naturally generated by the well-known Giudice–Masiero mechanism in the gravity-mediated supersymmetry breaking scenario [22], it leads to severe fine-tuning problems in precisely predicting the mass of the $Z$ boson, after considering the Large Hadron Collider (LHC) Higgs discovery and non-observations of any DM and supersymmetric particle signals in experiments when the theory runs down from an ultraviolet high-energy scale to the electroweak scale [23, 24, 25].
@@
The dilemma faced by the MSSM inspired us to extend this model. The Next-to-Minimal Supersymmetric Standard Model with a $\mathbb{Z}_{3}$ symmetry ($\mathbb{Z}_{3}$-NMSSM) introduces a gauge singlet superfield $\hat{S}$ on top of the MSSM [26, 27]. In this framework, the $\mu$-parameter of the MSSM is dynamically generated once the scalar component of the superfield $\hat{S}$ acquires a vacuum expectation value (vev) below approximately 1 TeV. This places it naturally at the electroweak scale. The neutralino sector includes an additional fermionic partner of $\hat{S}$, known as the Singlino. Both the Bino-dominated (which is the case in most physical scenarios) and Singlino-dominated neutralinos are viable DM candidates [28, 29, 30, 5, 31, 32, 33, 34, 35, 36, 15, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Primarily, the Bino-dominated DM candidate differs from the MSSM given the former’s ability to co-annihilate with a Singlino-dominated neutralino in achieving the measured abundance. Nonetheless, this scenario is confined to an exceedingly limited parameter space defined by specific conditions: $|2\kappa\mu/\lambda|\simeq|M_{1}|$, moderately large $\lambda$ and $\kappa$, and $\mu$ values exceeding 300 GeV [5, 15]. This space predicts a significant mixing of the SM Higgs field with the singlet Higgs field and is thus disfavored by the Higgs property measurements at the LHC. The characteristics of Singlino-dominated DM are contingent upon the values of $\lambda$, $\tan\beta$, $m_{\tilde{\chi}_{1}^{0}}$, and $\mu_{\rm eff}\equiv\lambda v_{s}/\sqrt{2}$  [47], where $\lambda$ denotes the singlet–doublet Higgs Yukawa coupling within the superpotential and $m_{\tilde{\chi}_{1}^{0}}\simeq 2\kappa\mu/\lambda$ requires $12|\kappa|/\lambda<$ 1. Considering that both SI and SD DM-nucleon scattering cross sections scale with $\lambda^{4}$ in the regime of heavy singlet Higgs bosons and consequently, the LZ experiment generally sets an upper limit of $\lambda\lesssim 0.1$, the DM is likely to achieve the observed abundance through one of two main mechanisms [31]: co-annihilating with Higgsino-dominated electroweakinos or undergoing resonant annihilations facilitated by singlet-dominated CP-even or CP-odd Higgs bosons. The former scenario is only viable within a narrow parameter space characterized by $2|\kappa|/\lambda\simeq 1$, $\lambda<0.1$, and $\mu<400$ GeV, where $\kappa$ represents the self-coupling coefficient of the singlet Higgs field [47]. To achieve the measured density, the latter scenario requires $2|m_{\tilde{\chi}_{1}^{0}}|$ to be close to the scalar mass. Thus, all these scenarios correspond to an exceptionally constrained parameter space, resulting in a significant suppression of Bayesian evidence [^1] Bayesian evidence is a fundamental concept in Bayesian statistics [48]. It is typically denoted as $Z(D|M)\equiv\int{P(D|O(M,\Theta))P(\Theta|M)\prod d\Theta_{i}}$, where $P(\Theta|M)$ represents the prior probability density function of input parameters $\Theta=(\Theta_{1},\Theta_{2},\cdots)$ in model $M$, and $P(D|O(M,\Theta))\equiv\mathcal{L}(\Theta)$ is the likelihood function for observed values $O$. The Bayesian evidence considers both the theoretical predictions $O(M,\Theta)$ and the experimental data $D$, and its computation involves marginalizing the model’s parameters, which requires integrating the likelihood function and prior probability over all possible parameter values. In Bayesian statistics, evidence is used to compare different models and determine which one is more plausible for explaining the observed data. While a higher value of $Z$ suggests that the respective model is more likely to agree with the data, an extremely small $Z$ means that the theory requires its parameters to be fine-tuned for the same..
@@ 2 Theoretical preliminaries
$W_{\rm GNMSSM}=W_{\text{Yukawa}}+\lambda\hat{S}\hat{H_{u}}\cdot\hat{H_{d}}+\frac{\kappa}{3}\hat{S}^{3}+\mu\hat{H_{u}}\cdot\hat{H_{d}}+\frac{1}{2}\mu^{\prime}\hat{S}^{2}+\xi\hat{S},$ (2.1)
@@ 2.1 Higgs sector of GNMSSM
$-\mathcal{L}_{soft}=$ $\Bigg{[}\lambda A_{\lambda}SH_{u}\cdot H_{d}+\frac{1}{3}\kappa A_{\kappa}S^{3}+m_{3}^{2}H_{u}\cdot H_{d}+\frac{1}{2}{m_{S}^{\prime}}^{2}S^{2}+\xi^{\prime}S+h.c.\Bigg{]}$ (2.2)
@@
with $v=\sqrt{v_{u}^{2}+v_{d}^{2}}\simeq 246~{}\mathrm{GeV}$, and the three masses are determined in solving the conditional equations to minimize the scalar potential. The Higgs sector is then described by eleven free parameters:
$\tan\beta,~{}\lambda,~{}\kappa,~{}v_{s},~{}A_{\lambda},~{}A_{\kappa},~{}\mu,~{}\mu^{\prime},~{}m_{3}^{2},~{}{m_{\text{S}}^{\prime}}^{2},~{}\xi^{\prime},$ (2.4)
@@
${\cal M}^{2}_{S,11}$ $=$ $\frac{\lambda v_{s}(\sqrt{2}A_{\lambda}+\kappa v_{s}+\sqrt{2}\mu^{\prime})+2m_{3}^{2}}{\sin 2\beta}+\frac{1}{2}(2m_{Z}^{2}-\lambda^{2}v^{2})\sin^{2}2\beta,$ (2.5)
@@
${\cal M}^{2}_{S,33}$ $=$ $\frac{(A_{\lambda}+\mu^{\prime})\sin 2\beta}{2\sqrt{2}v_{s}}\lambda v^{2}+\frac{\kappa v_{s}}{\sqrt{2}}(A_{\kappa}+2\sqrt{2}\kappa v_{s}+3\mu^{\prime})-\frac{\mu}{\sqrt{2}v_{s}}\lambda v^{2}-\frac{\sqrt{2}}{v_{s}}\xi^{\prime}.$
@@
${\cal M}^{2}_{P,22}$ $=$ $\frac{(A_{\lambda}+2\sqrt{2}\kappa v_{s}+\mu^{\prime})\sin 2\beta}{2\sqrt{2}v_{s}}\lambda v^{2}-\frac{\kappa v_{s}}{\sqrt{2}}(3A_{\kappa}+\mu^{\prime})$ (2.6)
@@
$\xi^{\prime}$ $=$ $\frac{v_{s}}{\sqrt{2}}\left[\frac{(A_{\lambda}+\mu^{\prime})\sin 2\beta}{2\sqrt{2}v_{s}}\lambda v^{2}+\frac{\kappa v_{s}}{\sqrt{2}}(A_{\kappa}+2\sqrt{2}\kappa v_{s}+3\mu^{\prime})-\frac{\mu}{\sqrt{2}v_{s}}\lambda v^{2}-m_{B}^{2}\right],$
$m_{S}^{\prime 2}$ $=$ $\frac{1}{2}\left[m_{B}^{2}-m^{2}_{C}+\lambda\kappa\sin 2\beta v^{2}-2\sqrt{2}\kappa v_{s}(A_{\kappa}+\frac{\kappa}{\sqrt{2}}v_{s}+\mu^{\prime})\right],$ (2.9)
@@
These formulas indicate that the Higgs mass matrices are determined by just eight out of the eleven parameters: $\tan\beta$, $\lambda$, $A_{\lambda}$, $m_{A}$, $m_{B}$, $m_{C}$, $m_{N}$, and $\mu_{tot}$. As will be shown in Eqs. (2.19-2.21), the remaining three parameters, $\kappa$, $A_{\kappa}$, and $v_{s}$, specifically influence triple Higgs coupling strengths.
@@ 2.3 DM physics of GNMSSM
In the GNMSSM, the following interactions are pertinent to the DM physics [26]:
@@
$C_{h_{i}h_{s}h_{s}}$ $=$ $\lambda v\,V_{h_{i}}^{\rm SM}V_{h_{s}}^{\rm S}V_{h_{s}}^{\rm S}(\lambda-\kappa\sin 2\beta)-\lambda\kappa vV_{h_{i}}^{\rm NSM}V_{h_{s}}^{\rm S}V_{h_{s}}^{\rm S}\cos 2\beta$ (2.19)
$+\sqrt{2}\kappa V_{h_{i}}^{\rm S}V_{h_{s}}^{\rm S}V_{h_{s}}^{\rm S}(3m_{N}+A_{\kappa})+C^{\prime}_{h_{s}}(\lambda,\kappa,\tan\beta,v_{s},A_{\lambda},m_{N}),$
$C_{h_{i}A_{s}A_{s}}$ $=$ $\lambda v\,V_{h_{i}}^{\rm SM}(\lambda+\kappa\sin 2\beta)+\lambda\kappa vV_{h_{i}}^{\rm NSM}\cos 2\beta$ (2.20)
$+\sqrt{2}\kappa V_{h_{i}}^{S}(m_{N}-A_{\kappa})+C^{\prime}_{A_{s}}(\lambda,\kappa,\tan\beta,v_{s},A_{\lambda},m_{N}),$
where the last contributions in $C_{h_{i}h_{s}h_{s}}$ and $C_{h_{i}A_{s}A_{s}}$ are suppressed by the Higgs mixings and they are functions of $\lambda$, $\kappa$, $\tan\beta$, $v_{s}$, $A_{\lambda}$, and $m_{N}$. In the small $\lambda$ case, we conclude that
$C_{\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}h_{s}}\simeq C_{\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}A_{s}}\simeq-\sqrt{2}\kappa,\quad C_{h_{s}h_{s}h_{s}}\simeq\sqrt{2}\kappa(3m_{N}+A_{\kappa}),$
$C_{h_{s}A_{s}A_{s}}\simeq\sqrt{2}\kappa(m_{N}-A_{\kappa}),$ (2.21)
and the other interactions, such as $C_{\tilde{\chi}^{0}_{1}\tilde{\chi}^{0}_{1}h}$, $C_{\tilde{\chi}^{0}_{1}\tilde{\chi}^{0}_{1}Z}$, $C_{hh_{s}h_{s}}$, and $C_{hA_{s}A_{s}}$, are suppressed by $\lambda$. These characteristics are crucial to understand the results of this study.
@@ 2.3.1 DM relic density
    This formula together with Eq. (2.21) indicate that $\left\langle\sigma v\right\rangle_{x_{F}}^{h_{s}h_{s}}$ depends mainly on $\kappa$, $A_{\kappa}$, $m_{h_{s}}$, and $m_{\tilde{\chi}_{1}^{0}}$. In particular, $\kappa$ plays a critical role in determining the magnitude of $\left\langle\sigma v\right\rangle_{x_{F}}^{h_{s}h_{s}}$, since most of its dominant contributions are proportional to $\kappa^{4}$. If $C_{h_{s}h_{s}h_{s}}$ is sufficiently small so that the s-channel contribution can be neglected, we have
@@
    respectively, assuming that $h_{s}$ is not exceptionally light to forbid $A_{s}$-mediated resonant annihilation. Furthermore, from the formula of $C_{h_{s}A_{s}A_{s}}$ in Eq. (2.21), one can infer that if $|m_{\tilde{\chi}_{1}^{0}}-A_{\kappa}|\ll(4m_{\tilde{\chi}_{1}^{0}}^{2}-m_{A_{s}}^{2})/|m_{\tilde{\chi}_{1}^{0}}|$, $|{\cal{A}}_{t}|$ is much larger than $|{\cal{A}}_{s}|$, reflecting that the $t$-channel contribution prevails. In this case, $\left\langle\sigma v\right\rangle_{x_{F}}^{h_{s}A_{s}}$ is simplified as
@@
    if $|m_{\tilde{\chi}_{1}^{0}}|$ is much larger than $m_{A_{s}}$ and $m_{h_{s}}$, and